OpenStudy (anonymous):

Write an appropriate direct variation equation if y=27 when x=9

OpenStudy (rockinhood):

We just went over these :D When doing direct variation y / x = k SO, 27 divided by 9 is 3. k = 3 Inverse Variation is you want to multiply to find each one ( xy = k )

OpenStudy (anonymous):

so the answer would only be k=3??

OpenStudy (rockinhood):

No, it'd be 27 / 9 = 3 using the y / x = k.

OpenStudy (rockinhood):

Sorry.